# Sharp systolic inequalities for invariant tight contact forms on principal $\mathbb{S}^{1}$-bundles over $\mathbb{S}^{2}$

### v2

@@ 1.1 The systolic ratio of a contact form
Motivated by the Riemannian setting, the authors of [38] defined the systole of a contact form $\alpha$ on a manifold $M$, which we write $\text{sys}(\alpha)$, as the minimal period of a periodic orbit of the Reeb flow of $\alpha$. If $M$ is a closed orientable manifold, this number is finite since Taubes proved in [42] that Reeb flows on compact three-manifolds have a periodic orbit. The sytolic ratio of $\alpha$ is defined as the ratio $\rho(\alpha)=\frac{(\text{sys}\alpha)^{2}}{\text{Vol}(\alpha)}$, where $\text{Vol}(\alpha)$ is the contact volume of $M$. Note that if $\alpha$ is the contact form induced on a unit cotangent bundle by a Riemannian metric, the metric and contact notions of systolic ratio coincide up to a dimensional factor.
@@ 1.2 Invariant tight contact forms and the main result
The classification of invariant contact structures on principal $\mathbb{S}^{1}$-bundles have been achieved in [35] and [27], see also [25, Section 5] for a detailed discussion of $\mathbb{S}^{1}$-invariant contact structures on $\mathbb{S}^{3}$. We give below examples of tight $\mathbb{S}^{1}$-invariant contact forms on bundles over $\mathbb{S}^{2}$.
  • The standard contact form $\alpha_{0}=\frac{1}{4i}\sum\limits_{i=1}^{2}\left(\bar{z_{i}}\mathrm{d}z_{i}-z_{i}\mathrm{d}\bar{z_{i}}\right)$ on $\mathbb{S}^{3}\subset\mathbb{C}^{2}$ is invariant under the $\mathbb{S}^{1}$-action $t\cdot(z_{1},z_{2})\mapsto(e^{2\pi it}z_{1},e^{2\pi it}z_{2})$ given by its Reeb flow. More generally, any positive Zoll contact form on a manifold $M$ is invariant under the $\mathbb{S}^{1}$-action induced by its Reeb flow, and the induced principal bundle structure on $M$ has negative Euler number. Zoll contact forms on non-trivial $\mathbb{S}^{1}$-principal bundles over a symplectic manifold with integral symplectic form can be obtained by the Boothby–Wang construction, see Section 7.2 in [24].
@@
    Riemannian metrics are special cases of Finsler metrics, where $F$ is the Euclidean norm induced by the metric. As in the Riemannian setting, a Finsler metric induces a contact form $\alpha_{F}$ on its unit tangent bundle. We call a Finsler metric on $\mathbb{S}^{2}$ a metric of revolution if $F$ is invariant under a circle action on $\mathbb{S}^{2}$ that has exactly two fixed points and is free on their complement. This action lifts to a free $\mathbb{S}^{1}$-action on $T^{1}M$ the unit tangent bundle for $F$. The induced principle bundle structure induced on $T^{1}M$ has Euler number 2. The contact form $\alpha_{F}$ is invariant under that action. Hence, $\alpha_{F}$ is in $\Omega(2)$.
@@
    Moreover, the supremum of $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ on $\Omega(e)$ is $\frac{1}{2}$ (and is not achieved).
@@
  • The fact that the supremum of $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ on $\Omega(e)$ is not achieved when $e>2$ can be understood as follows [^1] The author wishes to thank Christian Lange for pointing out that the systolic inequality had to be different for Euler numbers greater than 2.: the classification of Zoll contact forms in dimension 3, achieved in [14], implies that there is no positive Zoll contact form on $M$ when $e>2$. Moreover, contact forms that are local maximizers of the systolic ratio $\frac{\text{sys}(\alpha)^{2}}{\text{Vol}(\alpha)}$ in the class of $\mathbb{S}^{1}$-invariant contact forms are necessarily Zoll, as proved in [38, Theorem 3.4]. Thus the equality cannot be reached for the systolic inequality for invariant forms on $M$ when $e>2$.
@@ 1.3.2 Contractible systole and Besse contact forms
Theorem 1.2 also implies a systolic inequality for the contractible systole, stated in Corollary 1.5 below, whose proof will be given in 4.2. It recovers as a particular case Theorem A in [34]
@@ 1.3.3 Viterbo conjecture for invariant domains
Another corollary of Theorem 1.2 is an application to a conjecture of Viterbo. Let $\omega_{0}=\sum\limits_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$ be the standard symplectic form on $\mathbb{R}^{2n}$. We write $\text{Vol}(U,\omega_{0}^{n})$ for the symplectic volume of a set $U$. Viterbo stated the following conjecture in [43] (see for example [29] for an overview of the variations on this conjecture):
Conjecture 1.6 (Viterbo conjecture).
For any normalized symplectic capacity $c$ on $(\mathbb{R}^{2n},\omega_{0})$, and any bounded convex domain $U$ of $\mathbb{R}^{2n}$, we have
$c(U)^{n}\leq\text{Vol}(U,\omega_{0}^{n}).$
Moreover, equality holds if and only if $U$ is symplectomorphic to a ball.
Theorem 1.2 allows us to show that this conjecture holds in explicit situations, some of which are new.
@@
with equality if and only if $\alpha_{U}$ is Zoll, which by Proposition 4.3 in [6] is equivalent to the fact that $U$ is symplectomorphic to a ball.
For many normalized capacities $c$ it is known that $c(U)=\text{sys}(\alpha_{U})$ whenever $U$ is a smooth bounded convex domain. This is for example the case for the capacities $c_{HZ},c_{SH},c_{1}^{EH},c_{1}^{CH},c_{1}^{Alt}$, following the work of Ekeland, Hofer, Zehnder, Abbondandolo, Kang, Irie and Hutchings ([30] provides a related discussion). Following a consolidated tradition, we denote their common value on such domains by $c_{EHZ}$. As a consequence, Theorem 1.2 can be seen as a special case of Viterbo conjecture in dimension 4 for $c_{EHZ}$. This leads us to the corollary below.
@@
Identifying each $(x_{i},y_{i})$-plane of $\mathbb{R}^{4}$ with a copy of $\mathbb{C}$ yields two symplectic actions on $\mathbb{R}^{4}$. Moreover, those actions are free on $\mathbb{R}^{4}\setminus\{0\}$. Hence, as implied by Corollary 1.7, any convex domain $U$ that is invariant under any of those action satisfies the Viterbo conjecture. We have respectively $e=-1$ and $e=1$ for the induced actions of $\Phi^{+}$ and $\Phi^{-}$ on $\partial U$. As we will see from Proposition 2.4, this makes Corollary 1.7 essentially trivial for $\Phi^{+}$, but non-trivial and new for $\Phi^{-}$. The non-triviality of the result for $\Phi^{-}$ can be illustrated by the fact that the restriction of $\Phi^{-}$ to the unit sphere in $\mathbb{C}^{2}$ is a lift to $\mathbb{S}^{3}$ of the circle action on $T_{1}\mathbb{S}^{2}=\mathbb{R}P^{3}$ described in 1.3.1, which follows from the classification of $\mathbb{S}^{1}$-principal bundles over $\mathbb{S}^{2}$ by the Euler number. As such, Corollary 1.4 can be seen as a consequence of Corollary 1.7 for the action $\Phi^{-}$.
@@ 4 Proof of the main results
    $q(J^{\alpha}(r)-r(J^{\alpha})^{\prime}(r))$ $=q(f(|r|-rf^{\prime}(|r|)))$
    $\geq f(|r|-rf^{\prime}(|r|))$
@@
The contact form $\alpha$ constructed in the proof of Proposition 4.1 is briefly described in [34, Section 3], as a deformation of the round metric on $\mathbb{S}^{2}$ to a spindle orbifold.
@@
holds for similar reasons to (3). Furthermore, plugging the lower bound 6 in the equation (2) gives
@@
Equality holds if and only if the lifted form $p^{*}\alpha$ is Zoll. Since the contractible systolic ratio $\frac{\text{sys}_{\text{contr}}(\alpha)^{2}}{\text{Vol}(\alpha)}$ does not depend on $p$, whether or not $p^{*}\alpha$ is Zoll does not depend on the precise choice of covering map $p$. If $p^{*}\alpha$ is Zoll, then $\alpha$ is Besse since the Reeb orbits of $p^{*}\alpha$, which are closed, project on $M$ to the Reeb orbits of $\alpha$. However not all Besse forms on $M$ can be lifted to Zoll forms on $\mathbb{S}^{3}$. We spend the rest of the proof on proving a characterization of such Besse forms on $M$.
  • When $e<0$: if $p^{*}\alpha$ is Zoll, it follows for Proposition 2.4 that the lifted circle action on $\mathbb{S}^{3}$ coincides with the Reeb flow of $p^{*}\alpha$. Therefore the free circle action on $M$ coincides with the Reeb flow of $\alpha$, which is hence Zoll. Conversely, the Reeb flow of a Zoll contact form in $\Omega(e)$ coincides with the $\mathbb{S}^{1}$-action, which lifts to a free circle action on $\mathbb{S}^{3}$ via any degree $-e$ covering map. Hence Zoll contact forms in $\Omega(e)$ lift to Zoll contact forms on $\mathbb{S}^{3}$.
  • When $e>0$: if $p^{*}\alpha$ is Zoll, the proof of Theorem 1.2 implies that $K^{p^{*}\alpha}$ has exactly two critical points, meaning that exactly two Reeb orbits of $p^{*}\alpha$ coincide with orbits of the circle action. If $T>0$ is the minimal period of the Reeb flow of $p^{*}\alpha$, those two orbits correspond to two closed Reeb orbits of $\alpha$ of minimal period $\frac{T}{e}$. All the other orbits of $p^{*}\alpha$ project to closed Reeb orbits of $\alpha$ of period $T$ when $e$ is odd, $\frac{T}{2}$ when $e$ is even. Thus if $e\in\{1,2\}$, $\alpha$ is Zoll as well, but otherwise $\alpha$ is only Besse and has two singular orbits, which have the same minimal period. Conversely, Zoll contact forms in $\Omega(e)$ lift to Zoll contact forms on $\mathbb{S}^{3}$ when $e\in\{1,2\}$. Assume now that $e>2$. We claim that if $\alpha$ is a Besse contact form on $M$ with exactly two singular orbits, which have the same minimal period, then $\alpha$ lifts to a Zoll contact form on $\mathbb{S}^{3}$.
    The Reeb flow of any Besse contact form induces a Seifert fibration of $M$, namely a circle action on $M$ with finite isotropy groups. For any coprime integers $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$, we define a $\mathbb{S}^{1}$-action on $M$, whose orbits define a Seifert fibration of $M$, as follows : the $\mathbb{S}^{1}$-action on $\mathbb{S}^{3}$ defined by
@@
    As a consequence of the classification of Seifert fibrations on lens spaces (see [23, Theorem 5.1]), there exist coprime integers $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$ and a constant $T>0$ such that the Reeb flow of $\frac{\alpha}{T}$ is conjugated via a diffeomorphism of $M$ to the $\mathbb{S}^{1}$-action on $M$ associated to $(k_{1},k_{2})$. For the two singular orbits to have the same minimal period, $|k_{1}|$ and $|k_{2}|$ must be equal. Since $k_{1}$ and $k_{2}$ are coprime, they must belong to $\{-1,1\}$. If $k_{1}=-k_{2}$, the induced action on $M$ is free, hence $\alpha$ is Zoll. The proof of Theorem 1.2 implies that this cannot happen when $e>2$. Consequently $k_{1}=k_{2}$, and in that case the prime action spectrum of $\alpha$ is $\left(\frac{T}{e},T\right)$ if $e$ is odd, $\left(\frac{T}{e},\frac{T}{2}\right)$ if $e$ is even. It follows that $\frac{\alpha}{T}$ has the same prime action spectrum as $\alpha_{1}$. Applying Theorem 4.3, we get that there is a diffeomorphism $\psi$ of $M$ such that $\psi^{*}\alpha=\alpha_{1}$. Finally, this gives that $(\psi\circ p_{0})^{*}\alpha=p_{0}^{*}\alpha_{1}$ is a Zoll contact form on $\mathbb{S}^{3}$. As a consequence, when $e>2$ a Besse contact form on $M$ lifts to a Zoll contact form on $\mathbb{S}^{3}$ if and only if it has two singular orbits, which have the same prime period.

### v3

@@ 1.1 The systolic ratio of a contact form
Motivated by the Riemannian setting, the authors of [40] defined the systole of a contact form $\alpha$ on a manifold $M$, which we write $\text{sys}(\alpha)$, as the minimal period of a periodic orbit of the Reeb flow of $\alpha$. If $M$ is a closed orientable manifold, this number is finite since Taubes proved in [44] that Reeb flows on compact three-manifolds have a periodic orbit. The sytolic ratio of a contact form $\alpha$ on a contact manifold of dimension $12n+$ 1 is defined as the ratio $\rho(\alpha)=\frac{(\text{sys}\alpha)^{n+1}}{\text{Vol}(\alpha)}$, where $\text{Vol}(\alpha)$ is the contact volume of $M$. Note that if $\alpha$ is the contact form induced on a unit cotangent bundle by a Riemannian metric, the metric and contact notions of systolic ratio coincide up to a dimensional factor.
@@ 1.2 Invariant tight contact forms and the main result
The classification of invariant contact structures on principal $\mathbb{S}^{1}$-bundles have been achieved in [37] and [28], see also [26, Section 5] for a detailed discussion of $\mathbb{S}^{1}$-invariant contact structures on $\mathbb{S}^{3}$. We give below examples of tight $\mathbb{S}^{1}$-invariant contact forms on bundles over $\mathbb{S}^{2}$.
  • The standard contact form $\alpha_{0}=\frac{1}{4i}\sum\limits_{i=1}^{2}\left(\bar{z_{i}}\mathrm{d}z_{i}-z_{i}\mathrm{d}\bar{z_{i}}\right)$ on $\mathbb{S}^{3}\subset\mathbb{C}^{2}$ is invariant under the $\mathbb{S}^{1}$-action $t\cdot(z_{1},z_{2})\mapsto(e^{2\pi it}z_{1},e^{2\pi it}z_{2})$ given by its Reeb flow. The corresponding $\mathbb{S}^{1}$-bundle has Euler number $-1$. More generally, any positive Zoll contact form on a manifold $M$ is invariant under the $\mathbb{S}^{1}$-action induced by its Reeb flow, and the induced principal bundle structure on $M$ has negative Euler number. Zoll contact forms on non-trivial $\mathbb{S}^{1}$-principal bundles over a symplectic manifold with integral symplectic form can be obtained by the Boothby–Wang construction, see Section 7.2 in [25].
@@
    Riemannian metrics are special cases of Finsler metrics, where $F$ is the Euclidean norm induced by the metric. As in the Riemannian setting, a Finsler metric induces a contact form $\alpha_{F}$ on its unit tangent bundle. We call a Finsler metric on $\mathbb{S}^{2}$ a metric of revolution if $F$ is invariant under a circle action on $\mathbb{S}^{2}$ that has exactly two fixed points and is free on their complement. This action lifts to a free $\mathbb{S}^{1}$-action on $T^{1}\mathbb{S}^{2}$ the unit tangent bundle for $F$. The induced principle bundle structure induced on $T^{1}\mathbb{S}^{2}$ has Euler number 2. The contact form $\alpha_{F}$ is invariant under that action. Hence, $\alpha_{F}$ is in $\Omega(2)$.
@@
    Moreover, the supremum of the systolic ratio $\rho(\alpha)$ on $\Omega(e)$ is $\frac{1}{2}$ (and is not achieved).
@@
  • The fact that the supremum of $\rho(\alpha)$ on $\Omega(e)$ is not achieved when $e>2$ can be understood as follows [^1] The author wishes to thank Christian Lange for pointing out that the systolic inequality had to be different for Euler numbers greater than 2.: the classification of Zoll contact forms in dimension 3, achieved in [15], implies that there is no positive Zoll contact form on $M$ when $e>2$. Moreover, contact forms that are local maximizers of the systolic ratio $\rho(\alpha)$ in the class of $\mathbb{S}^{1}$-invariant contact forms are necessarily Zoll, as proved in [40, Theorem 3.4]. Thus the equality cannot be reached for the systolic inequality for invariant forms on $M$ when $e>2$.
@@ 1.3.2 Contractible systole and Besse contact forms
Theorem 1.2 also implies a systolic inequality for the contractible systole, stated in Corollary 1.5 below, whose proof will be given in 4.2. It recovers as a particular case Theorem A in [36]. Depending on the Euler number of the circle action, the maximizers of the contractible systolic ratio are either Zoll contact forms, or Besse contact forms with exactly two singular orbits, which have the same minimal period. When $e>2$, Besse contact forms with two singular orbits of same minimal period can be obtained as a quotient of the standard contact form $\alpha_{0}$ on $\mathbb{S}^{3}\subset\mathbb{C}^{2}$ via the $\mathbb{Z}/e\mathbb{Z}$ action generated by $(z_{1},z_{2})\mapsto(e^{\frac{2\pi i}{e}}z_{1},e^{-\frac{2\pi i}{e}}z_{2})$. If $T$ is the minimal period of the Reeb flow $\alpha_{0}$, the contact form induced on the quotient has two singular orbits of period $\frac{T}{e}$. The other Reeb orbits have period $T$ if $e$ is odd, $\frac{T}{2}$ if $e$ is even. Alternatively, those Besse contact forms can be described as spindle orbifolds as in [36].
@@ 1.3.3 Viterbo conjecture for invariant domains
Another corollary of Theorem 1.2 is an application to a conjecture of Viterbo. Let $\omega_{0}=\sum\limits_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$ be the standard symplectic form on $\mathbb{R}^{2n}$. We write $\text{Vol}(U,\omega_{0}^{n})$ for the symplectic volume of a set $U\subset\mathbb{R}^{2n}$. In [45] (see also [30] for an overview of variations of this conjecture), Viterbo conjectures that for any symplectic capacity $c$ and any open bounded convex domain $U\subset\mathbb{R}^{2n}$ the inequality
$c(U)^{n}\leq\text{Vol}(U,\omega_{0}^{n})$ (1)
holds, and that equality holds if and only if $U$ is symplectomorphic to a ball. He also proves the inequality up to a dimensional factor. Since any convex domain $U$ in $\mathbb{R}^{2n}$ is of contact type, its boundary carries a contact form $\alpha_{U}$. For many normalized capacities $c$ it is known that $c(U)=\text{sys}(\alpha_{U})$ whenever $U$ is a smooth bounded convex domain. This is for example the case for the capacities $c_{HZ},c_{SH},c_{1}^{EH},c_{1}^{CH},c_{1}^{Alt}$, following the work of Ekeland, Hofer, Zehnder, Abbondandolo, Kang, Irie and Hutchings ([32] provides a related discussion). By a consolidated tradition, we denote their common value on such domains by $c_{EHZ}$. As a consequence, systolic inequalities for convex domains can be seen as special cases of Viterbo conjecture.
It is now known that Viterbo conjecture does not hold at this level of generality. Indeed, in [31], the authors constructed explicit examples of bounded convex domains $U\subset\mathbb{R}^{4}$ satisfying $\frac{c_{EHZ}(U)^{2}}{\text{Vol}(U,\omega_{0}^{2})}=\frac{\sqrt{5}+3}{5}>1$, as well as examples in any higher dimension, disproving Viterbo conjecture. On the other hand, Viterbo conjecture has been confirmed for several subclasses of convex domains (see [31] and the references therein). Furthermore, in view of its connection with Mahler’s conjecture in convex geometry (see [11]), it is interesting to explore the range of its validity. Theorem 1.2 allows us to show that Viterbo conjecture holds in explicit situations, some of which are new.
@@
with equality if and only if $\alpha_{U}$ is Zoll, which by Proposition 4.3 in [6] is equivalent to the fact that $U$ is symplectomorphic to a ball. This leads us to the corollary below.
@@
Identifying each $(x_{i},y_{i})$-plane of $\mathbb{R}^{4}$ with a copy of $\mathbb{C}$ yields two symplectic actions on $\mathbb{R}^{4}$. Those two actions are not symplectomorphic, since they define two principal bundle structures on $\mathbb{S}^{3}$ with different Euler number, hence not conjugated via an orientation preserving diffeomorphism. Moreover, those actions are free on $\mathbb{R}^{4}\setminus\{0\}$. Hence, as implied by Corollary 1.6, any convex domain $U$ that is invariant under any of those action satisfies the Viterbo conjecture. We have respectively $e=-1$ and $e=1$ for the induced actions of $\Phi^{+}$ and $\Phi^{-}$ on $\partial U$. As we will see from Proposition 2.4, this makes Corollary 1.6 essentially trivial for $\Phi^{+}$, but non-trivial and new for $\Phi^{-}$. The non-triviality of the result for $\Phi^{-}$ can be illustrated by the fact that the restriction of $\Phi^{-}$ to the unit sphere in $\mathbb{C}^{2}$ is a lift to $\mathbb{S}^{3}$ of the circle action on $T_{1}\mathbb{S}^{2}=\mathbb{R}P^{3}$ described in 1.3.1, which follows from the classification of $\mathbb{S}^{1}$-principal bundles over $\mathbb{S}^{2}$ by the Euler number. As such, Corollary 1.4 can be seen as a consequence of Corollary 1.6 for the action $\Phi^{-}$.
@@ 4 Proof of the main results
    $q(J^{\alpha}(r)-r(J^{\alpha})^{\prime}(r))$ $=q(f(|r|)-rf^{\prime}(|r|))$
    $\geq f(|r|)-rf^{\prime}(|r|)$
@@
The contact form $\alpha$ constructed in the proof of Proposition 4.1 is briefly described in [36, Section 3], as a deformation of the round metric on $\mathbb{S}^{2}$ to a spindle orbifold.
@@
holds for similar reasons to (4). Furthermore, plugging the lower bound 7 in the equation (3) gives
@@
Equality holds if and only if the lifted form $p^{*}\alpha$ is Zoll. Since the contractible systolic ratio $\frac{\text{sys}_{\text{contr}}(\alpha)^{2}}{\text{Vol}(\alpha)}$ does not depend on $p$, whether or not $p^{*}\alpha$ is Zoll does not depend on the specific choice of covering map $p$. If $p^{*}\alpha$ is Zoll, then $\alpha$ is Besse since the Reeb orbits of $p^{*}\alpha$, which are closed, project on $M$ to the Reeb orbits of $\alpha$. However not all Besse forms on $M$ can be lifted to Zoll forms on $\mathbb{S}^{3}$. We spend the rest of the proof on proving a characterization of such Besse forms on $M$.
  • When $e<0$: if $p^{*}\alpha$ is Zoll, it follows from Proposition 2.4 that the lifted circle action on $\mathbb{S}^{3}$ coincides with the Reeb flow of $p^{*}\alpha$. Therefore the free circle action on $M$ coincides with the Reeb flow of $\alpha$, which is hence Zoll. Conversely, the Reeb flow of a Zoll contact form in $\Omega(e)$ coincides with the $\mathbb{S}^{1}$-action, which lifts to a free circle action on $\mathbb{S}^{3}$ via any degree $-e$ covering map. Hence Zoll contact forms in $\Omega(e)$ lift to Zoll contact forms on $\mathbb{S}^{3}$.
  • When $e>0$: if $p^{*}\alpha$ is Zoll, the proof of Theorem 1.2 implies that $K^{p^{*}\alpha}$ has exactly two critical points, meaning that exactly two Reeb orbits of $p^{*}\alpha$ coincide with orbits of the circle action. If $T>0$ is the minimal period of the Reeb flow of $p^{*}\alpha$, those two orbits correspond to two closed Reeb orbits of $\alpha$ of minimal period $\frac{T}{e}$. All the other orbits of $p^{*}\alpha$ project to closed Reeb orbits of $\alpha$ of period $T$ when $e$ is odd, $\frac{T}{2}$ when $e$ is even. Thus if $e\in\{1,2\}$, $\alpha$ is Zoll as well, but otherwise $\alpha$ is only Besse and has two singular orbits, which have the same minimal period.
    Conversely, Zoll contact forms in $\Omega(e)$ lift to Zoll contact forms on $\mathbb{S}^{3}$ when $e\in\{1,2\}$. Assume now that $e>2$. We claim that if $\alpha$ is a Besse contact form on $M$ with exactly two singular orbits, which have the same minimal period, then $\alpha$ lifts to a Zoll contact form on $\mathbb{S}^{3}$.
    Indeed, the Reeb flow of any Besse contact form induces a Seifert fibration of $M$, namely a circle action on $M$ with finite isotropy groups. For any coprime integers $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$, we define a $\mathbb{S}^{1}$-action on $M$, whose orbits define a Seifert fibration of $M$, as follows : the $\mathbb{S}^{1}$-action on $\mathbb{S}^{3}$ defined by
@@
    As a consequence of the classification of Seifert fibrations on lens spaces (see [24, Theorem 5.1]), there exist coprime integers $k_{1},k_{2}\in\mathbb{Z}\setminus\{0\}$ and a constant $T>0$ such that the Reeb flow of $\frac{\alpha}{T}$ is conjugated via a diffeomorphism of $M$ to the $\mathbb{S}^{1}$-action on $M$ associated to $(k_{1},k_{2})$. For the two singular orbits to have the same minimal period, $|k_{1}|$ and $|k_{2}|$ must be equal. Since $k_{1}$ and $k_{2}$ are coprime, they must belong to $\{-1,1\}$. If $k_{1}=-k_{2}$, the induced action on $M$ is free, hence $\alpha$ is Zoll. The proof of Theorem 1.2 implies that this cannot happen when $e>2$. Consequently $k_{1}=k_{2}$, and in that case the prime action spectrum of $\alpha$ is $\left(\frac{T}{e},T\right)$ if $e$ is odd, $\left(\frac{T}{e},\frac{T}{2}\right)$ if $e$ is even. It follows that $\frac{\alpha}{T}$ has the same prime action spectrum as $\alpha_{1}$. Applying Theorem 4.3, we get that there is a diffeomorphism $\psi$ of $M$ such that $\psi^{*}\alpha=\alpha_{1}$. Finally, this gives that $(\psi\circ p_{0})^{*}\alpha=p_{0}^{*}\alpha_{1}$ is a Zoll contact form on $\mathbb{S}^{3}$. As a consequence, when $e>2$ a Besse contact form on $M$ lifts to a Zoll contact form on $\mathbb{S}^{3}$ if and only if it has two singular orbits, which have the same prime period. This conclude the proof.